OpenStudy (anonymous):

How do I set this p? Use the compound interest formula to compute the total amount accumulated. $10,000 for 20 years at 5% compounded annually. A. $20,000.00 B. $25,269.50 C. $19,500.00 D. $26,532.98

OpenStudy (anonymous):

post the formula

OpenStudy (anonymous):

A=P(1+r/n)^(nt)

OpenStudy (anonymous):

\[ P= P_0\left( 1 + r) \right)^n \]

OpenStudy (anonymous):

r as a decimal rather than a percentage, n as the number of compoundings per year, and t as the time in years.

OpenStudy (anonymous):

\[ 10000(1.05)^{20} \]

OpenStudy (anonymous):

Looks right

OpenStudy (anonymous):

. $26,532.98

OpenStudy (anonymous):

10,000 *(1+.05)^(20) D

OpenStudy (anonymous):

that s 2 diferent formulas you guys are givin me. so witch one is it?

OpenStudy (anonymous):

Yep. That looks right. @LilRyG - They're both right

OpenStudy (anonymous):

well the one with the n the n is just one. :)

OpenStudy (anonymous):

Except n means something different in her formula.

OpenStudy (anonymous):

n represents the number of years. Hers is only good for annual compounding.

OpenStudy (anonymous):

A=P(1+r/n)^(nt) the n is used to manipulate it to yearly quarterly or monthly

OpenStudy (anonymous):

with 1 it is yearly

OpenStudy (anonymous):

with n=12 montly n=12/4 quarterly

OpenStudy (anonymous):

okay thank you guys!!

OpenStudy (anonymous):

de nada :)